13% of all Americans live in poverty. If 49 Americans are randomly selected, find the probability that

a. Exactly 8 of them live in poverty. _____
b. At most 4 of them live in poverty. ______
c. At least 3 of them live in poverty. _____
d. Between 2 and 7 (including 2 and 7) of them live in poverty. _____
Answer #1

Sal! Given, P = 13.1. - 0.13 n=49 8 a) 49 49-8 (1-0.13) P(x=8) (0.13) 2 Cg = 0.1919 (6) P( x4) {p(x=4) 4) +PCX= 3) + P(x = 2)